OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

x=1+4y Substituting in the 2nd eqn gives 2(1+4y)-9y=3 2+8y-9y=3 -y=1 y=-1 x=1-4 =-3 (-3,-1)